OpenStudy (anonymous):

Evaluate: 5[1 9 2 -6]

jimthompson5910 (jim_thompson5910):

I'm assuming this is a matrix

jimthompson5910 (jim_thompson5910):

if so, then you multiply the 5 by each number inside the matrix

jimthompson5910 (jim_thompson5910):

Example: \[\Large 7*\begin{bmatrix} 8 & 5\\ 2 & -3 \end{bmatrix} = \begin{bmatrix} 7*8 & 7*5\\ 7*2 & 7*(-3) \end{bmatrix} = \begin{bmatrix} 56 & 35\\ 14 & -21 \end{bmatrix} \] This is NOT the answer to your question as it's just an example

OpenStudy (anonymous):

so: 5*1=5 5*9=45 5*2=10 5*(-6)=-30

OpenStudy (anonymous):

Then the answer would be: [5 45 10 -30] ??? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that looks good

OpenStudy (anonymous):

thank you for helping me understand the concepts of matrix for multiplication

jimthompson5910 (jim_thompson5910):

you're welcome this is really called "scalar multiplication" or "multiplying a matrix by a scalar" because the 5 outside is the scalar (a single number) the matrix is the collection of numbers (the nonscalar)

jimthompson5910 (jim_thompson5910):

matrix multiplication is some other concept you'll get to